Packaging Solutions Corporation manufactures and sells a wide variety of packaging products. Performance reports are prepared monthly for each department. The planning budget and flexible budget for the Production Department are based on the following formulas, where q is the number of labor-hours worked in a month:

Cost Formulas
Direct labor $16.10q
Indirect labor $4,300 + $1.70q
Utilities $5,400 + $0.80q
Supplies $1,300 + $0.20q
Equipment depreciation $18,600 + $2.50q
Factory rent $8,100
Property taxes $2,600
Factory administration $13,600 + $0.60q

The Production Department planned to work 4,200 labor-hours in March; however, it actually worked 4,000 labor-hours during the month. Its actual costs incurred in March are listed below:

Actual Cost Incurred in March
Direct labor $ 65,940
Indirect labor $ 10,620
Utilities $ 9,150
Supplies $ 2,350
Equipment depreciation $ 28,600
Factory rent $ 8,500
Property taxes $ 2,600
Factory administration $ 15,350

Required:

1. Prepare the Production Department’s planning budget for the month.

2. Prepare the Production Department’s flexible budget for the month.

3. Prepare the Production Department’s flexible budget performance report for March, including both the spending and activity variances.
Answer #1

(1). Planning budget will be made with q of 4200:

Cost formula Planning budget when q = 4200
Direct labor $16.10q 67,620.00
Indirect labor $4,300 + $1.70q 11,440.00
Utilities $5,400 + $0.80q 8,760.00
Supplies $1,300 + $0.20q 2,140.00
Equipment depreciation $18,600 + $2.50q 29,100.00
Factory rent $8,100 8,100.00
Property taxes $2,600 2,600.00
Factory administration $13,600 + $0.60q 16,120.00

Calculations: In the above table the values are computed with q as 4200. Thus direct labor = 16.10*4200 = 67,620. Indirect labor = 4300 + (1.7*4200) = 11,440. Similarly other values have been computed.

(2): Flexible budget will be made with a q of 4,000.

Cost formula Flexible budget when q = 4000
Direct labor $16.10q 64,400.00
Indirect labor $4,300 + $1.70q 11,100.00
Utilities $5,400 + $0.80q 8,600.00
Supplies $1,300 + $0.20q 2,100.00
Equipment depreciation $18,600 + $2.50q 28,600.00
Factory rent $8,100 8,100.00
Property taxes $2,600 2,600.00
Factory administration $13,600 + $0.60q 16,000.00

(3): Spending variance = actual figures - flexible budget figures

Actual results Flexible budget Spending variance Favorbale/Unfavorable
Direct labor 65,940.00 64,400.00 1,540.00 U
Indirect labor 10,620.00 11,100.00 480.00 F
Utilities 9,150.00 8,600.00 550.00 U
Supplies 2,350.00 2,100.00 250.00 U
Equipment depreciation 28,600.00 28,600.00 0.00
Factory rent 8,500.00 8,100.00 400.00 U
Property taxes 2,600.00 2,600.00 0.00
Factory administration 15,350.00 16,000.00 650.00 F

When actual expenses > expenses as per flexible budget then the spending variance is unfavorable (U) and when actual expenses < flexible amounts then variance is favorable (F)

Activity variance = flexible budget - planning budget. When flexible expenses < planned expenses then the activity variance is favorable, otherwise it is unfavorable.

Planning budget Flexible budget Activity variance Favorbale/Unfavorable
Direct labor 67,620.00 64,400.00 3,220.00 F
Indirect labor 11,440.00 11,100.00 340.00 F
Utilities 8,760.00 8,600.00 160.00 F
Supplies 2,140.00 2,100.00 40.00 F
Equipment depreciation 29,100.00 28,600.00 500.00 F
Factory rent 8,100.00 8,100.00 0.00
Property taxes 2,600.00 2,600.00 0.00
Factory administration 16,120.00 16,000.00 120.00 F